write the equation in standard form
9(x-4)^2+ 16(y-3)^2=144

Answer by nyc_function(2741)

Divide all terms by 144. The goal is to get 1 on the right side of the equation.
9(x-4)^2/144 + 16(y-3)^2/144 = 144/144
[(x - 4)^2]/16 + [(y - 3)^2]/9 = 1